A cylindrical can is to be constructed to hold a fixed volume of liquid. The cost of the


material used for the top and bottom of the can is 3 cents per square inch, and the cost of the


material used for the curved side is 2 cents per square inch. Use calculus to derive a simple


relationship between the radius and height of the can that is the least costly to construct.

1
Expert's answer
2021-11-30T14:47:40-0500

Solution;

"Volume=\u03c0r^2h"

Cost of material;

"C=3(2\u03c0r^2)+2(2\u03c0rh)=6\u03c0r^2+4\u03c0rh"

"But ;"

"h=\frac{V}{πr^2}"

Substituting into the cost;

"C=6πr^2+4πr(\frac{V}{πr^2})"

"C=6πr^2+4\frac Vr"

Differentiate cost w.r.t r;

"\frac{dC}{dr}=12πr-\frac{4V}{r^2}=0"

We have ;

"r^3=\frac{4V}{12π}=\frac{V}{3π}"

"r=(\frac{V}{3π})^{\frac13}"

While;

"h=\frac{V}{πr^2}=\frac{V}{π(\frac{V}{3π})^{\frac13}}"

Hence;

"h=2.08(\frac{V}π)^{\frac13}=1.42V^{\frac13}"